Please help!
4th sqrt of 48 plus 4th sqrt of 3 equals what?

Answer by ankor@dixie-net.com(22740)

4th sqrt of 48 plus 4th sqrt of 3 equals what?
:
4thRoot(48) + 4thRoot(3) =
:
Factor to reveal the 4th power of 2 which is 16
4thRoot(16*3) + 4thRoot(3) =
:
Extract the 4th root of 16
2*4thRoot(3) + 4thRoot(3) = 3*4thRoot(3)
:
These are like terms, so just add them
:
"
Check using a calc; enter: 48^(1/4) + 3^(1/4) = 3.94382..
Enter our solution 3 * 3^(1/4) = 3.94382...; confirms our solution